OpenStudy (anonymous):

If R = (2a,0), T = (2c,2d), and S = (2b,0) and M and N are midpoints of RT and ST what is the length of MN?

OpenStudy (anonymous):

M = (2a +2c)/2,2d/2 = (a+c,d) N = (2b + 2c)/2, 2d/2 = (b+c,d) \[L = \sqrt{((b+c)-(a+c))^2 + (d-d)^2}\] \[L = \sqrt{(b-a)^2}\] or \[L = \left| b-a \right|\]